# Observation of non-Hermitian dislocation bound states and dislocation skin effects

### v1

@@ Abstract
As illustrated in Fig. 2(a), two distinct midgap states are clearly observed in the ${\rm M}$ phase, with their spatial profiles localized at the dislocation cores, see Fig. 2(b). The dislocation modes are separated from the bulk states, see Fig. 2(c). By contrast, the $\Gamma$ phase reveals only bulk eigenstates, with no midgap energies detected, see Figs. 2(d) and 2(e). These findings are consistent with the theoretical prediction that a Hermitian Chern insulator hosting a dislocation supports zero-energy states only in the ${\rm M}$ phase, where the first Chern number $C=-1$ [4], see also Figs. 2(f)- 2(j).
@@
Upon introducing moderate NH perturbations, the energy spectra become considerably more intricate, as shown in Fig. 3. See SM for corresponding theoretical predictions [56]. Nevertheless, two NHDS remain pinned near zero-energy within the line gap between 1038 Hz and 1042 Hz (approximately), distinguishing them from the point-gap associated with typical NHSE in other NH lattices [34]. The spatial distribution of the corresponding eigenstates reveals their pronounced dependence on the NH perturbations in the system. Specifically, Fig. 3(b) shows that when only $h_{x}$ is nonzero, the weight of the right eigenstates is biased along the $x$-direction, with a greater concentration on the right side ($+x$) of the dislocation cores. This observation is consistent with the fact that this NH perturbation breaks (preserves) the reflection symmetry about $y$-axis ($x$-axis). Similarly, Fig. 3(e) depicts that for only a nonzero $h_{y}$, the eigenstate weight shifts along the $y$-direction, with an enhanced intensity near the bottom dislocation core, as this NH perturbation breaks (preserves) the reflection symmetry about $x$-axis ($y$-axis). By contrast, when only $h_{z}$ is nonzero, the intrinsic $C_{4}$ symmetry of $H_{\mathrm{H}}(\mathbf{k})$ is preserved, resulting in an approximately symmetric eigenstate distribution along both $x$ and $y$ directions, see Fig. 3(h).
@@
Furthermore, the geometry of dislocations (encoded in the Burgers vector) plays a crucial role in determining the fate of the D-NHSE under PBCs. As shown in Fig. 1, the dislocation core possesses one (two) nearest-neighbor sites in the $y$ ($x$) direction when the Burgers vector is $\mathbf{b}=\pm\mathbf{e}_{x}$, thus mimicking the coordination number of a site living on an $y$ directional boundary. Consequently, under PBC, no NHSE is observed for nonzero $h_{x}$ near the defect core, see Fig. 3(c), while nonzero $h_{y}$ yields a D-NHSE as shwon in Fig. 3(f). However, D-NHSE is relatively weak compared to the conventional NHSE observed under OBC at the exterior boudanries of the system. See the SM for a detailed comparison [56]. Notably, when only $h_{z}$ is finite, no NHSE is observed anywhere in the system, irrespective of the boundary conditions.
Figure 4: Similar to Fig. 3, but for strong NH perturbations, showing the absence of NHDS. Complex energy spectrum for (a) $\mathbf{h}=(3.3\,\mathrm{Hz},0,0)$ (b) $\mathbf{h}=(0,3.3\,\mathrm{Hz},0)$, and (c) $\mathbf{h}=(0,0,3.3\,\mathrm{Hz})$. Corresponding total amplitude of all the right eigenstates are shown in (d), (e), and (f), respectively.
@@
Melting of NHDS. As the strength of NH perturbations increases, initially well-localized NHDS gradually spread away from the defect cores. Under PBC, where the conventional NHSE is not supported, the weight of these states is progressively absorbed into the bulk. This trend is observed with different perturbation levels $h_{j}/t_{0}=0.3,0.6,0.9$ for $j=x,y,z$, as detailed in the SM [56]. A similar behavior is found for the left eigenstates. Nonetheless, in the presence of dislocations, the NH Chern insulators with $C=-1$ continues to support topologically robust localized states around the defect cores, which happens for $|h_{j}|<t_{0}$ when $|m_{0}|=t_{0}$.
@@
For the parameter setting $t_{0}=-m_{0}=3\,\mathrm{Hz}$, the EPs, where the real and imaginary parts of the eigenvalues of $H_{\mathrm{NH}}(\mathbf{k})$ vanish, first appear in the Brillouin zone when $h_{j}=3\,\mathrm{Hz}$ for $j=x,y,z$. As the system approaches such NH band gap closing, NHDS begin to lose their localization near the defect cores and gradually merge into the bulk states. This melting process is demonstrated in Fig. 4, where the system is placed slightly above the critical NH perturbation $h_{j}=3.3\,\mathrm{Hz}$ for $j=x,y,z$, further detailed in the SM. In the strong NH perturbation regime, the NHDS near zero-energy disappear, and only the summed weight of all right eigenstates is presented in Fig. 4(d)-(f), showing that the localized defect states have fully merged into the bulk under PBC. As further corroborated by experiments under OBC (see SM), the NHDS are found to completely merge into the skin modes living at the exterior boundaries of the system. Collectively, Figs. 34, and the SM provide clear evidence for a localization-delocalization phase transition in the NHDS, a phenomenon exclusively observable in NH systems.
@@
It is noteworthy that Fig. 4(e) illustrates a pronounced D-NHSE in the absence of NHDS for stronger $h_{y}$, with a higher right eigenstate weight at the top dislocation core compared to that for a weaker $h_{y}$, shown in Fig. 3(f). By contrast, for nonzero $h_{x}$ and $h_{z}$, the D-NHSE is either absent (likelihood) or much weaker, as shown in Figs. 4(d) and 4(f), respectively. Therefore, our experimental obseravtion clearly demonstrates D-NHSE that is distinct from the conventional point-gap supported NHSE under OBC [34] and depends on the relative orientation of the Burgers vector and the directionality of the NHSE.
Figure 5: Experimental observations of the total amplitude of all right eigenstates under OBCs in the ${\rm M}$ phase with $t_{0}=-m_{0}=3\,\mathrm{Hz}$ for (a) $\mathbf{h}=(0,0,0)$ (Hermitian limit), (b) $\mathbf{h}=(1.8\,\mathrm{Hz},0,0)$, (c) $\mathbf{h}=(0,1.8\,\mathrm{Hz},0)$, and (d) $\mathbf{h}=(0,0,1.8\,\mathrm{Hz})$.
@@
Observations under OBC. Under OBC, the spatial distribution of the summed weight of all right eigenstates reveals valuable insights. For a Hermitian insulator and a NH one with nonzero $h_{z}$, such a weight extends uniformly throughout the entire lattice, as shown in Figs. 5(a) and 5(d), respectively, confirming the absence of any NHSE therein. However, with the NH perturbations $h_{x}$ and $h_{y}$ a clear signature of NHSE emerges, with the right wavefunctions localizing toward the positive $x$ and $y$ boundaries, respectively, shown in Figs. 5(b) and 5(c). This anisotropic behavior stems from the broken reflection and inversion symmetries by the NH perturbations $h_{x}$ and $h_{y}$. With the emergence of NHSE along the edges, the dislocation cores themselves do not exhibit a pronounced NHSE under OBC for nonzero $h_{x}$ and $h_{y}$. By contrast, when the Burgers vector is perpendicular to the direction of the NHSE observed under OBC, a D-NHSE becomes visible under PBC, as show in Fig. 4(e).
@@
Summary and discussions. To summarize, our experiments provide conclusive evidences for NHDS and D-NHSE in a 2D acoustic Chern insulator. In the Hermitian limit, dislocations host zero-energy states in the ${\rm M}$ phase, consistent with the bulk-defect correspondence [1, 4, 5]. Under moderate NH perturbations, these states persist within the line gap. As the perturbation strength increases, we observe a localization-delocalization transition for NHDS as they gradually spread and merge with bulk states under PBCs when the system approaches the shore of hosting EPs in the Brillouin zone [38]. By contrast, the D-NHSE appears for any arbitrary strength of NH perturbations, when the direction of the conventional NHSE is orthogonal to the Burgers vector. Our results not only bridge the gap between NH band topology and defect physics but also establish crystalline defects as a universal tool for probing NH topological matter.
@@
Acknowledgments. Y.J. thanks the support of startup funds from Pennsylvania State University and NSF awards 2039463, 195122, and 2401236. B.R. was supported by NSF CAREER Grant No. DMR-2238679.

### v2

@@ 2.5 Observations of NHDS and D-NHSE
As illustrated in Fig. 4A, two distinct midgap states are clearly observed only in the $\rm M$} phase, with their spatial profiles localized at the dislocation cores, see Fig. 4B. The dislocation modes are separated from the bulk states, see Fig. 4C. By contrast, the $\Gamma$ phase reveals only bulk eigenstates, with no midgap energies detected, see Figs. 4D1 and 4E1. These findings are consistent with the theoretical prediction that a Hermitian Chern insulator hosting a dislocation supports zero-energy states only in the $\rm M$} phase, where the first Chern number C=1C=-1 [4], see also Figs. 4D2 and 4E2.
@@
Upon introducing moderate NH perturbations, the energy spectra become considerably more intricate, as shown in Fig. 5. Nevertheless, two NHDS remain pinned near zero-energy within the line gap between 1038 Hz and 1042 Hz (approximately), distinguishing them from the point-gap associated with typical NHSE in other NH lattices [34]. The spatial distribution of the corresponding eigenstates reveals their pronounced dependence on the NH perturbations in the system. Specifically, Figs. 5A2 and 5B2 show that when only $h_{x}$ is nonzero, the weight of the right eigenstates is biased along the $x$-direction, with a greater concentration on the right side (+x+x) of the dislocation cores. This observation is consistent with the fact that this NH perturbation breaks (preserves) the reflection symmetry about $y$-axis ($x$-axis). Similarly, Figs. 5C2 and 5D2 depict that for only a nonzero $h_{y}$, the eigenstate weight shifts along the $y$-direction, with an enhanced intensity near the bottom dislocation core, as this NH perturbation breaks (preserves) the reflection symmetry about $x$-axis ($y$-axis). By contrast, when only $h_{z}$ is nonzero, the intrinsic $C_{4}$ symmetry of $H_{\mathrm{H}}(\mathbf{k})$ is preserved, resulting in an approximately symmetric eigenstate distribution along both $x$ and $y$ directions, see Figs. 5E2 and 5F2.
@@
Furthermore, the geometry of dislocations (encoded in the Burgers vector) plays a crucial role in determining the fate of the D-NHSE under PBCs. As shown in Fig. 1A, the dislocation core possesses one (two) nearest-neighbor sites in the $y$ ($x$) direction when the Burgers vector is $\mathbf{b}=\pm\mathbf{e}_{x}$, thus mimicking the coordination number of a site living on an $y$ directional boundary. Consequently, under PBCs, no NHSE is observed for nonzero $h_{x}$ near the defect cores, see Figs. 5A3 and 5B3 and SI Appendix Figs. S24–S39, while nonzero $h_{y}$ yields a D-NHSE as shown in Figs. 5C3 and 5D3. However, D-NHSE is relatively weak compared to the conventional NHSE observed under OBCs at the exterior boundaries of the system. See SI Appendix Figs. S40–S55 for a detailed comparison. Notably, when only $h_{z}$ is nonzero, no NHSE is observed anywhere in the system, irrespective of the boundary conditions, see Figs. 5E3 and 5F3 and SI Appendix Figs. S56–S71.
2.6 Observations of the NHDS melting
@@
As the strength of NH perturbations increases, initially well-localized NHDS gradually spread away from the defect cores. Under PBCs, where the conventional NHSE at the exterior boundaries of the system is not supported, the weight of these states is progressively absorbed into the bulk. This trend is observed with different perturbation levels $h_{j}/t_{0}=0.3,0.6,0.9$ for $j=x,y,z$, as detailed in SI Appendix Figs. S24–S35, S40–S51, and S56–S67. A similar behavior is found for the left eigenstates. Nonetheless, in the presence of dislocations, the NH Chern insulators with C=1C=-1 continues to support topologically robust localized states around the defect cores, which happens for $|h_{j}|<t_{0}$ when $|m_{0}|=t_{0}$.
@@
For the parameter setting $t_{0}=-m_{0}=3\,\mathrm{Hz}$, the EPs first appear in the Brillouin zone when $h_{j}=3\,\mathrm{Hz}$ for $j=x,y,z$, as validated in Fig. 3. As the system approaches such a NH band gap closing, NHDS begin to lose their localization near the defect cores and gradually merge into the bulk states. This melting process is demonstrated in Fig. 6, where the system is placed slightly above the critical NH perturbation $h_{j}=3.3\,\mathrm{Hz}$ for $j=x,y,z$, further detailed in SI Appendix Figs. S39–S39, S52–S55, and S68–S71. In the strong NH perturbation regime, the NHDS near zero-energy disappear, and only the summed weight of all right eigenstates is presented in Figs. 6A2–F2, showing that the localized defect states have fully merged into the bulk under PBC. As further corroborated by experiments under OBCs (see SI Appendix Figs. S37, S53, and S69), the NHDS are found to completely merge into the skin states living at the exterior boundaries of the system. Collectively, Figs. 56, and the SI Appendix Figs. S24–S71, provide clear evidence for a localization-delocalization phase transition in the NHDS, a phenomenon exclusively observable in NH systems.
@@
It is noteworthy that Figs. 6B2 and 6E2 illustrate a pronounced D-NHSE in the absence of NHDS for stronger $h_{y}$, with a higher right eigenstate weight at the top dislocation core compared to that for a weaker $h_{y}$, shown in Figs. 5C3 and 5D3. By contrast, for nonzero $h_{x}$ and $h_{z}$, the D-NHSE is either absent (likelihood) or much weaker, as shown in Figs. 6A2 and 6C2, respectively. Therefore, our experimental observations clearly demonstrate a D-NHSE that is distinct from the conventional point-gap-supported NHSE under OBCs [34], and it depends on the relative orientation of the Burgers vector and the directionality of the NHSE.
2.7 Observations under OBCs
@@
Under OBCs, the spatial distribution of the summed weight of all right eigenstates reveals valuable insights. For a Hermitian insulator and a NH one with nonzero $h_{z}$, such a weight extends uniformly throughout the entire lattice, as shown in Figs. 7A and 7D, respectively, confirming the absence of any NHSE therein. However, with the NH perturbations $h_{x}$ and $h_{y}$, clear signatures of NHSE emerge, with the right wavefunctions localizing toward the positive $x$ and $y$ boundaries, respectively, shown in Figs. 7B and 7C. This anisotropic behavior stems from the broken reflection and inversion symmetries by the NH perturbations $h_{x}$ and $h_{y}$. With the emergence of NHSE along the edges, the dislocation cores themselves do not exhibit a pronounced NHSE under OBCs for nonzero $h_{x}$ and $h_{y}$. By contrast, when the Burgers vector is perpendicular to the direction of the NHSE observed under OBCs, a D-NHSE becomes visible under PBC, as show in Figs. 6B2 and 6E2.
3 Conclusion
@@
To summarize, our experiments provide conclusive evidences for NHDS and D-NHSE in a 2D acoustic Chern insulator. In the Hermitian limit, dislocations host zero-energy states in the $\rm M$} phase, consistent with the bulk-defect correspondence [1, 4, 5]. Under moderate NH perturbations, these states persist within the line gap. As the perturbation strength increases, we observe a localization-delocalization transition for NHDS as they gradually spread and merge with bulk states under PBCs when the system approaches the shore of hosting EPs in the Brillouin zone [38]. By contrast, the D-NHSE appears for any arbitrary strength of NH perturbations, when the direction of the conventional NHSE is orthogonal to the Burgers vector.
Beyond their fundamental significance, these phenomena point toward specific applications in acoustic device engineering. For instance, the NHDS, which exhibits line-gap protected localization only within a specific spectral window, can serve as the basis for frequency-selective acoustic concentrators or filters. Conversely, the D-NHSE, which funnels bulk energy toward the dislocation core, suggests a mechanism for high-sensitivity sensing. The resulting intensity accumulation at the defect core can naturally amplify the system’s response to minute local perturbations, such as mass loading or structural defects. Our results not only bridge the gap between NH band topology and defect physics but also establish crystalline defects as a universal tool for probing NH topological matter.
4 Experimental Section
[…]
Experimental setup
Figure 2A shows a photograph of our experimental setup. The system consists of 56 3D-printed acoustic cavities, each representing a site in the NH lattice. A custom-designed digital controller is employed to precisely tune the onsite potentials of the acoustic cavities and the hoppings between them. The tuning of onsite potentials and hoppings is implemented using active components, specifically microphone-loudspeaker (detector-source) pairs. The hopping implementations are independent of spatial distances between sites, enabling flexible implementation of various boundary conditions and lattice geometries with ease.
The custom-made digital controller comprises three components: the core board, the motherboard, and the input/output (IO) board. The core board houses a field-programmable gate array (FPGA, XC7K325T, Xilinx) and a digital signal processor (DSP, TMS320C6678, Texas Instruments), which enable real-time signal processing. During experiments, the system operates at a sampling frequency of 12.8 kHz. The motherboard supplies power and facilitates communication between the core and IO boards. The IO board handles analog signal acquisition and output, interfacing with microphones and loudspeakers, respectively. Each IO board supports 16 channels for both analog inputs and outputs. The input channels utilize analog-to-digital converters (ADCs, ADC7606B, ADI), and the output channels employ digital-to-analog converters (DACs, DAC8568, Texas Instruments).
[…]
Acoustic cavities and tuning of onsite potentials
SI Appendix Sec. S2.2 shows the tuning process of the onsite potentials of acoustic cavities. The acoustic cavities used in this study were fabricated using 3D printing with a tolerance of 0.2 mm or within 0.3%. The material is LEDO 6060 photosensitive resin, which behaves as acoustically rigid for airborne sound. Each printed cavity has a wall thickness of 6 mm. As shown in Fig. 2B, the cavities are hexagonal prisms with an interior height of $l=164\,\mathrm{mm}$ and a side length of $25\,\mathrm{mm}$. To characterize the cavity in the experiments, a loudspeaker (source) excites the cavity, and a microphone (detector) measures the acoustic pressure. The onsite potential, $\omega_{0}$, of a cavity is retrieved using the Green’s function for a single site:
[…]
where $\omega$ is the excitation frequency.
[…]
Tuning of hoppings
Figure 2D shows the tuning process of the hoppings. Both nonreciprocal and reciprocal hoppings between the cavities are implemented using detector-source (microphone-loudspeaker) pairs. In our platform, the hopping strength and phase are precisely controlled by a digital multi-channel controller, allowing flexible and reconfigurable manipulation of lattice hoppings. Each unidirectional hopping is realized through a loudspeaker, a microphone, and an audio amplifier (Texas Instruments LM386). The loudspeaker and microphone are positioned at the bottom of each cavity. The microphone captures the acoustic pressure signal, which is processed by the controller to adjust its phase and amplitude before being emitted by the loudspeaker in the connected cavity. The analytical solution of this model is described in SI Appendix Sec. S2.4. Crucially, to avoid self-oscillation, we actively verify that the system remains in the stable regime by confirming that all eigenenergies in the reconstructed complex spectrum possess negative imaginary parts (see SI Appendix Sec. 2.6 for details).
The tuning of the hopping parameters is performed as follows. As illustrated in Fig. 2D, a microphone is placed inside cavity 1 to detect the acoustic pressure. The detected pressure signal is then phase-adjusted, amplified, and emitted by the loudspeaker positioned in cavity 2. To determine both the amplitude and the phase of the unidirectional hopping, we excite at cavity 1 using a loudspeaker (measurement source) and measure the acoustic pressure in both cavities using two microphones (measurement detector). The cross-power spectral density of these two measured pressure signals is expressed by
$P(\omega)\equiv\frac{\psi_{\mathrm{2}}(\omega)}{\psi_{\mathrm{1}}(\omega)}=\frac{\kappa_{0}}{\omega-\omega_{0}},$ (4)
where $\kappa_{0}$ is the tuned hopping.
[…]
Implementation of boundary conditions
Leveraging the fully programmable nature of our active coupling scheme, we can seamlessly switch between PBCs and OBCs without physical reconfiguration of the acoustic structure. Since all hoppings are implemented via microphone-loudspeaker pairs connected through digital circuits, PBCs are realized by electronically enabling the feedback loops that connect the boundary sites across the lattice edges. Conversely, OBCs are achieved simply by disabling these specific boundary coupling channels in the controller. A schematic illustration using a simplified one-dimensional chain model is provided in SI Appendix Sec. S2.7 to clarify this implementation.
[…]
Measurement process
To obtain the full Green’s function matrix, an acoustic source (loudspeaker) is sequentially excited at each cavity site, and the acoustic pressure is measured in all cavities using microphones. This procedure is systematically repeated for every site across the entire lattice. The complex-valued energy spectra as well as left and right eigenstates of NH acoustic Chern insulators are obtained based on the method proposed in Ref. [49].
Supporting Information
Supporting Information is available from the Wiley Online Library or from the author.
Note added.
During the review process of this manuscript, we became aware of a related work by Wu et al. [57]. We note that the two works explore distinct and complementary regimes of non-Hermitian defect physics. Wu et al. investigate the point-gap topology in a 2D Hatano-Nelson lattice, focusing on the D-NHSE. Conversely, our work realizes a line-gap non-Hermitian Chern insulator, which enables the observation of the coexistence of D-NHSE and topological defect-bound states (NHDS), as well as the interplay with EPs. Collectively, these studies establish a comprehensive experimental framework for understanding defect engineering across different non-Hermitian topological classes.
Acknowledgements
@@
Y.J. thanks the support of startup funds from Pennsylvania State University and NSF awards 2039463, 195122, and 2401236. B.R. was supported by NSF CAREER Grant No. DMR-2238679.